OpenStudy (anonymous):

THIS IS A MATH QUESTION!!! 55. During one year, people in the United States older than 18 years old watched a total of 342 billion hours of television. The population of the United States older than 18 years old was about 209 million people. Part A: On average, how many hours of television did each person older than 18 years old watch that year? Round to the nearest hour. Part B: On average, how many hours per week did each person older than 18 years old watch that year? Round to the nearest hour.

OpenStudy (anonymous):

Oh thank god someone is here!!

OpenStudy (anonymous):

I was beginning to think that NOBODY was gonna help me!!!

OpenStudy (anonymous):

oh, hi!!! thank you for coming here!!! I really dont understand how to work this problem!!! I mean, once I have the equation, I know how to work it but I cant seem to make an equation!!

OpenStudy (anonymous):

ok, so the answer in my book is about 31 hours, how would I get to that?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I added part A, but I already have the answer for it!!! the answer to part A is about 1636 hours!!!

OpenStudy (anonymous):

is that a 0 or a 9 above the 10?

OpenStudy (anonymous):

my bad:\[\frac{ 342 \times 10^9 }{ 52 } \times \frac{ 1 }{ 209 \times 10^6 } = 31.47\]

OpenStudy (anonymous):

a 9

OpenStudy (anonymous):

oh thank you so much!!! thank you for helping me!! I will HAVE to fan you now!!! :D THANK YOU!!!

OpenStudy (anonymous):

lol c: happy to help